OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 3y = 6 x = -3/2y + 3 3y = -2x + 6 y = -2/3x + 2

OpenStudy (anonymous):

the last one

OpenStudy (akshay_budhkar):

the 4 th option as the form is y=mx+C

OpenStudy (anonymous):

slope intercept is y = mx + b....so just follow that

OpenStudy (angela210793):

subtract 2x both sides 3y=6-2x divide 3 both sides y=6/3-2x/3 y=2-2x/3